# Volatility Parametrizations with Random Coefficients: Analytic Flexibility for Implied Volatility Surfaces

### v2

@@ Abstract
It is a market practice to express market-implied volatilities in some parametric form. The most popular parametrizations are based on or inspired by an underlying stochastic model, like the Heston model (SVI method [9]) or the SABR model (SABR-parametrization [15]). Their popularity is often driven by a closed-form representation enabling efficient calibration. However, these representations indirectly impose a model-specific volatility structure on observable market quotes. When the market’s volatility does not follow the parametric model regime, the calibration procedure will fail or lead to extreme parameters, indicating inconsistency. This article addresses this critical limitation - we propose an arbitrage-free framework for letting the parameters from the parametric implied volatility formula be random. The method enhances the existing parametrizations and enables a significant widening of the spectrum of permissible shapes of implied volatilities while preserving analyticity and, therefore, computation efficiency. We demonstrate the effectiveness of the novel method on real data from short-term index and equity options, where the standard parametrizations fail to capture market dynamics. Our results show that the proposed method is particularly powerful in modeling the implied volatility curves of short expiry options preceding an earnings announcement, when the risk-neutral probability density function exhibits a bimodal form.
@@ 1 Introduction
The existing literature on this problem is elaborate, and more innovative techniques are continuously being developed. This paper aims to provide a generic method to enable implied volatility surface parametrizations to become more effective. Inspired by the RAnD method [12], we will demonstrate a technique to improve the flexibility of the parametrizations by randomization one of the parameters to obtain a new parametrization with only limited additional parameters. The additional flexibility helps the parametrization to accurately model the market volatility when the classical parametrizations fail, such as before earnings announcements or generally for short-maturity options.
@@ 1.1 Problem Setting
Secondly, the resulting volatility surface should at all times be free of static arbitrage opportunities. With the Black-Scholes equation, the volatility surface defines the theoretical prices for put and call options on the entire domain. If these option prices present arbitrage opportunities by buying or selling options with different strikes and expiries, the volatility surface is not deemed suitable since market participants could exploit these, thereby eliminating the opportunities. The absence of arbitrage is guaranteed by a number of conditions that can be equivalently formulated on the option prices or directly on the volatility surface [14].
Once the implied volatility function $\hat{\sigma}(T,K)$ is constructed from the market quotes and is arbitrage-free, it has various applications. Firstly, the option surface provides information on the market’s expectation of the probability distribution of underlying asset $S$. For instance, the risk-neutral probability distribution of the asset $S$ can be derived with the Breeden-Litzenberger formula [3], and the surface can be used to calibrate pricing models for exotic derivatives. Furthermore, the surface can be used for trading purposes by assessing volatility expectations, detecting arbitrage opportunities, or providing liquidity to the market.
@@ 1.2 Literature Review
Implied volatility surface modeling has been studied extensively in the past. Successful approaches have been developed using stochastic and local volatility models, statistical properties of the implied volatility surface, interpolation schemes, and, lately, machine learning. An extensive review of the methods is found in [16]. We will quickly summarize the approaches that are relevant to our methodology.
The simplest way to obtain the function $\hat{\sigma}(T,K)$ is to choose an interpolation and extrapolation scheme to interpolate the available volatility quotes on the continuum. It is well known [7] that such an approach is not arbitrage-free since a direct interpolation on the volatility surface often leads to a non-convex option pricing function. The mispricing is further exaggerated for assets with few option quotes available in the market, where the interpolation determines a large portion of the surface, increasing the risk of generating implausible or arbitrageable prices. An arbitrage-free interpolation scheme is feasible on the option pricing function rather than the implied volatilities [18, 2]. The interpolation methods remain problematic for assets for which only a few high-quality option prices are available. Furthermore, to obtain the implied volatilities from the market prices, an inversion of the Black-Scholes formula is required, which can only be achieved by a root-finding algorithm.
@@
Volatility parametrizations are often directly derived from a parametric asset model, such as a stochastic volatility model. Although a volatility surface can be generally derived for any asset price model, not all models offer a convenient analytical form for the volatility surface. Defining an analytically tractable and realistic model (in terms of fitting the market) is thus a challenging task. The Stochastic Volatility Inspired (SVI) parametrization [9] was derived from a stochastic volatility model and enjoyed great success in the past. The parametrization was extended and improved in multiple works [19, 6]. Another successful parametrization is the SABR parametrization [15], a direct result of the Hagan et al. formula provided by the SABR model (Stochastic-Alpha-Beta-Rho). Parametric volatility surfaces are convenient for numerous reasons. Firstly, since the parametrizations are derived from an asset model, the absence of arbitrage is guaranteed as long as the parameters remain in a well-defined range. Simple constraints on the parameter space are usually sufficient to ensure the absence of arbitrage. Secondly, the parameters can be attributed to the shape of the volatility surface. This allows for a simple way to compare surfaces and the expression of the dynamics of the surface in terms of the dynamics of the parameter. Thirdly, theoretically, no minimum number of market quotes is required to fit an arbitrage-free surface. Although a higher number of quotes is desired for stability in the calibration, a parametrization can be fit to any number of quotes and remain arbitrage-free. This is of particular importance for assets with only a few quotes or only in a particular strike region, where little to no information is provided for the tails of the volatility curves.
@@
The main drawback of parametrizations is that there is no guarantee that a surface that is appropriately close to the market data can be found. The flexibility of the surface is limited by the chosen parametrization, which does not cover all viable (i.e. arbitrage-free) volatility shapes. If the market conditions are not within the scope of the parametrization, the fitting process fails, and the parametrization misrepresents the actual market conditions. These out-of-scope market conditions can be systemic, meaning that parametric models are generally unfit to model a certain market behavior. For instance, it is well-known that short-maturity options in equity markets exhibit a steeper at-the-money implied volatility term structure than attainable under regular stochastic volatility models [8]. Alternatively, the market condition can be out of scope due to an irregular behavior in the market, such as spikes in volatility or higher-than-usual uncertainty. It is not uncommon for very short-term options markets (near expiry) to exhibit a W-shaped volatility shape, or “mustache” shapes [10] before an earnings announcement of the underlying equity. These shapes of volatility surface arise from bimodal risk-neutral probability density functions for the stock price, which reflect the dichotomous nature of the earnings, which can either have a positive or negative effect on the asset price. Classical diffusion models [^1] It needs to be mentioned that SVI-type parametrizations do not rely on the an underlying model, but their development was merely inspired by volatility shapes of the Heston model. Nevertheless, one observes the same characteristics in terms of fit as classical stochastic volatility models. are unable to create such shapes due to the single continuous diffusive driver, and the corresponding volatility parametrization is limited in such cases. This limitation highlights the industry’s need for more flexible parameterized surfaces to achieve better fits in such market situations.
@@ 1.3 Contributions
In this paper, we study volatility surfaces in cases when the parametric surfaces reach the limit of their flexibility. We introduce a generic method to enhance the flexibility in such cases using a randomization scheme of the parameters of the surface. By replacing one or more parameters of the parametric surface with a random variable, we derive a new implied volatility surface with increased flexibility to fit the market. The new randomized volatility surface is expressed in terms of the original parameters plus parameters to specify the chosen random variable. To show the effectiveness of the novel method, we apply the method to the existing parametrizations of the SABR parametrization and demonstrate that the additional flexibility enables them to better fit the market data for SPX options with expiry dates of less than half a year. Furthermore, we derive a second kind of randomization where we randomize the spot price $S_{0}$ of the asset. This arbitrage-free randomization is shown to be particularly effective for near-maturity options (i.e., options with an expiry date in the next few days) before an earning announcement, which induces a multi-modal-type volatility regime. [^2] We speak of a multi-modal-type volatility regime when the risk-neutral probability density functions implied by the volatility surface exhibit more than one mode, and are therefore multi-modal. The paper is organized in the following way: Section 2 describes the process of randomizing a parametrization of volatility surfaces and proves that the randomization is indeed free of static arbitrage. We present a few examples of randomized parametrizations and their effectiveness in Section 3.2. In particular, we show an improved fit to options on the S&P 500 index. In Section 4, we then introduce the randomized spot parametrization and also present an example of its use on short-maturity AMZM options on the day of an earnings announcement. Finally, we conclude the paper in Section 5.
@@ 2.1 General Volatility Parametrizations
We consider a type of implied volatility surfaces, which are given in parametric form, defined by a set of parameters. As a starting point, suppose that $\overline{p}:=(p_{1},\dots,p_{m})$ is an array of $m$ constant parameters on a parameter domain $\overline{p}\in\mathcal{D}\subset\mathbb{R}^{m}$. We define a parameterized implied volatility function as a positive function
@@
for an asset with spot price $S_{0}$ at $t_{0}$. The function provides a continuum of European call and put option prices on the underlying asset on $(t_{0}\times\infty)\times[0,\infty)$ given the implied volatility $\hat{\sigma}(T,K;\overline{p})$. We can extend the functions to the limit points at $T=t_{0}$, which, under certain conditions, exist are given by $V_{c}(t_{0},K;\overline{p})=(S_{0}-K)^{+}$ and $V_{c}(t_{0},K;\overline{p})=(K-S_{0})^{+}$ (see 2.1). The pricing function defines two additional surfaces in $\mathbb{R}^{3}$ over the domain $\Pi$, which we refer to as the (put/call) pricing surface.
Since the parameterized implied volatility and pricing surface define a set of prices for tradable options, these prices must be free of arbitrage opportunities. There are various ways to define the absence of arbitrage of an implied volatility surface, most of which are equivalent. Here, we will utilize the model-free definition [21], which does not rely on the introduction of stochastic asset models but can be expressed as a set of conditions on the pricing surfaces.
@@
Equivalent conditions in terms of the put-call parity can be derived for the put pricing function $V_{p}(T,K;\overline{p})$, or directly in terms of the volatility surface $\hat{\sigma}(T,K;\overline{p})$ [14]. Under these conditions, one can prove the existence of a non-negative local martingale process on a suitable probability space such that the call price function can be written as a risk-neutral expectation of the final payoff. In other words, an arbitrage-free model (price process) exists that yields the volatility surface. However, we will not utilize this fact to avoid introducing any model dynamics and remain “model-free”.
Due to the continuity of the Black-Scholes formula with respect to the volatility input, and the conditions above, it is necessary for the parameterized volatility function and parameterized price function $V_{c/p}(T,K;\overline{p})$ to be $\mathcal{C}^{1,2}$, meaning that the price and volatility surfaces are both smooth. Therefore, one can obtain the risk-neutral probability density functions $\{p_{S_{t}}\colon t\geq t_{0}\}$, defined as [^4] Note that one can also use the put prices $V_{p}(t,x;\overline{p})$ to obtain the same result.
@@
for any $t\geq t_{0}$ using the Breeden-Litzenberger formula [3] and risk-free rate $r$. In particular, one can show with a few manipulations that expected value of the random variable $S_{t}$ whose PDF is given by is $p_{S_{t}}(\cdot)$, is
@@
Conversely, given a set of PDFs, such that (2.4) holds for all $t\geq t_{0}$, we can define an arbitrage-free pricing surface by integration of the probability densities [9].
@@
with respect to a desired norm $\lVert$.$\rVert$ . ∥. The ability to fit the market, therefore, directly depends on the “size” of $\mathcal{S}$, whether $\mathcal{S}$ contains a function that can match the market.
@@ 2.2 Volatility Parametrizations with Random Parameters
In an effort to increase the size of $\mathcal{S}$ and obtain better fitting arbitrage-free volatility surfaces, we consider the possibility of adding stochasticity to the implied volatility surface by replacing one of the parameters $p_{i}$ for $1\leq i\leq m$ with a random variable. Let $\hat{\sigma}(T,K;\overline{p})$ be a parameterized implied volatility function with parameter domain $\mathcal{D}\subset\mathbb{R}^{m}$ and suppose that $\vartheta$ is an absolutely continuous real-valued random variable on a probability space $(\Omega,\mathcal{F},\mathbb{P})$ with associated probability density function $f_{\vartheta}(\cdot)$. We can replace the $i$-th entry of $\overline{p}$ with the random variable $\vartheta$ to obtain the random vector $\overline{p}(\vartheta)=(p_{1},p_{2},\dots,\vartheta,\dots,p_{m})$. If we assume that $\overline{p}(\vartheta)$ is almost surely in the domain $\mathcal{D}$, a realization $\omega\in\Omega$ determines a real-valued vector $\overline{p}(\theta):=(p_{1},p_{2},\dots,\theta,\dots,p_{m})\in\mathbb{R}^{m}$ with $\theta=\vartheta(\omega)$, which almost surely provides an arbitrage-free implied volatility surface $\hat{\sigma}(T,K;\overline{p}(\theta))$ and its associated pricing functions $V_{c/p}(T,K;\overline{p}(\theta))$.
Assuming measurability of the parametrization function in the parameters, the pricing function $V_{c/p}(T,K;\overline{p}(\vartheta))$ is thus a random variable, and we can compute the expectation
@@
where $f_{\vartheta}$ is the associated probability density function of $\vartheta$. The equation shows that the expected option price under the randomized volatility surface is effectively an average of option prices on the sample space of the parameters of the volatility surface. Since the expectation of the randomized pricing function is a deterministic function in two variables, we can investigate its suitability to be a pricing function. Due to the linear properties of probability density functions, the function is indeed an arbitrage-free pricing function.
@@
Let $\hat{\sigma}(T,K;\overline{p})$ be an arbitrage-free parameterized implied volatility function on the parameter domain $\overline{p}\in\mathcal{D}$ and let $\vartheta$ be a real-valued absolutely continuous random variable. Suppose that $\overline{p}(\vartheta)=(p_{1},p_{2},\dots,\vartheta,\dots,p_{m})$ is the random vector where we replaced the $i$-th parameter of $\overline{p}$ with $\vartheta$. Then, the map $\hat{\sigma}(T,K)$ such that
@@
$V_{c/p}(T,K;\overline{p}(\theta))={\rm e}^{-r(T-t_{0})}\int_{K}^{\infty}(x-K)p_{S_{T};\theta}(x){\rm d}x.$
@@
$\mathbb{E}[V_{c}(T,K;\overline{p}(\vartheta))]$ $=\int_{\mathbb{R}}V_{c}(T,K;\overline{p}(\theta))f_{\vartheta}(\theta){\rm d}\theta$
$=\int_{\mathbb{R}}{\rm e}^{-r(T-t_{0})}\int_{K}^{\infty}(x-K)p_{S_{T};\theta}(x){\rm d}xf_{\vartheta}(\theta){\rm d}\theta$
$={\rm e}^{-r(T-t_{0})}\int_{K}^{\infty}(x-K)\int_{\mathbb{R}}p_{S_{T};\theta}(x)f_{\vartheta}(\theta){\rm d}\theta{\rm d}x.$
Since the random vector $\overline{p}(\vartheta)$ takes values almost surely in $\mathcal{D}$, the function $p_{S_{T};\theta}(\cdot)$ is a proper probability density function when $f_{\vartheta}(\theta)>0$. The expression $\int_{\mathbb{R}}p_{S_{T};\theta}(x)f_{\vartheta}(\theta){\rm d}\theta$ is a convex combination of probability densities (mixture density), which we define as the function
$\overline{f}(x):=\int_{\mathbb{R}}p_{S_{T};\theta}(x)f_{\vartheta}(\theta){\rm d}\theta,$
We claim $\overline{f}(x)$ to be a proper probability density function on $[0,\infty)$. The function $\overline{f}(x)$ is certainly positive since $p_{S_{T};\theta}(\cdot),f_{\vartheta}(\cdot)$ are positive functions, and integrates to 1 since
$\int_{0}^{\infty}\overline{f}(x){\rm d}x$ $=\int_{0}^{\infty}\int_{\mathbb{R}}p_{S_{T};\theta}(x)f_{\vartheta}(\theta){\rm d}\theta{\rm d}x$
$=\int_{\mathbb{R}}\int_{0}^{\infty}\ p_{S_{T};\theta}(x){\rm d}xf_{\vartheta}(\theta){\rm d}\theta$
$=\int_{\mathbb{R}}f_{\vartheta}(\theta){\rm d}\theta$
@@
This proves the claim and is a sufficient condition for $\hat{\sigma}_{T}(K)$ to be butterfly-arbitrage-free [9]. It remains to show that condition v) is true as well. Let $K$ be fixed and suppose that $t_{1}<t_{2}\in\Pi_{T}$. Since $\hat{\sigma}(T,K;\overline{p}(\theta))$ is arbitrage-free, we have that $V_{c}(t_{1},K;\overline{p}(\theta))\leq V_{c}(t_{2},K;\overline{p}(\theta))$ for $\overline{p}(\theta)\in\mathcal{D}$. Since this is almost surely the case, we conclude that

### v3

@@ Abstract
It is a market practice to express market-implied volatilities in some parametric form. The most popular parametrizations are based on or inspired by an underlying stochastic model, like the Heston model (SVI method [Quantitative Finance, 14 (2014), pp. 59-71]) or the SABR model (SABR-parametrization [Wilmott Magazine, (2002), pp. 84-108]). Their popularity is often driven by a closed-form representation enabling efficient calibration. However, these representations indirectly impose a model-specific volatility structure on observable market quotes. When the market’s volatility does not follow the parametric model regime, the calibration procedure will fail or lead to extreme parameters, indicating inconsistency. This article addresses this critical limitation - we propose an arbitrage-free framework for letting the parameters from the parametric implied volatility formula be random. The method enhances the existing parametrizations and enables a significant widening of the spectrum of permissible shapes of implied volatilities while preserving analyticity and, therefore, computation efficiency. We demonstrate the effectiveness of the novel method on real data from short-term index and equity options, where the standard parametrizations fail to capture market dynamics. Our results show that the proposed method is particularly powerful in modeling the implied volatility curves of short expiry options preceding an earnings announcement, when the risk-neutral probability density function exhibits a bimodal form.
@@ 1 Introduction
The existing literature on this problem is elaborate, and more innovative techniques are continuously being developed. This paper aims to provide a generic method to enable implied volatility surface parametrizations to become more effective by means of parameter randomization. It has been shown in the past [jacquier2019randomized, jacquier2018black, github], that randomization of model parameters in stochastic models can increase their flexibility and, therefore, enable a better fit to the market. We apply a randomization technique to improve the flexibility of implied volatility parametrizations and obtain a new parametrization with only limited additional parameters. The extra flexibility helps the parametrization to accurately model the market volatility when the classical parametrizations fail, such as before earnings announcements or generally for short-maturity options.
@@ 1.1 Problem Setting
Secondly, the resulting volatility surface should always be free of static arbitrage opportunities. With the Black-Scholes equation, the volatility surface defines the theoretical prices for put and call options on the entire domain. If these option prices present arbitrage opportunities by buying or selling options with different strikes and expiries, the volatility surface is not deemed suitable since market participants could exploit these, thereby eliminating the opportunities. The absence of arbitrage is guaranteed by several conditions that can be equivalently formulated on the option prices or directly on the volatility surface [guo2016generalized].
Once the implied volatility function $\hat{\sigma}(T,K)$ is constructed from the market quotes and is arbitrage-free, it has various applications. Firstly, the option surface provides information on the market’s expectation of the probability distribution of underlying asset $S$. For instance, the risk-neutral probability distribution of the asset $S$ can be derived with the Breeden-Litzenberger formula [breeden1978prices], and the surface can be used to calibrate pricing models for exotic derivatives. Furthermore, the surface can be used for trading purposes by assessing volatility expectations, detecting arbitrage opportunities, or providing liquidity to the market.
@@ 1.2 Literature Review
Implied volatility surface modeling has been studied extensively in the past. Successful approaches have been developed using stochastic and local volatility models, statistical properties of the implied volatility surface, and interpolation schemes. As of late, new streams of data-driven approaches have been considered, further enhanced with machine learning tools [franccois2022venturing, ackerer2020deep, gonon2024operator, chataigner2020deep]. An extensive review of the classical methods is found in [homescu2011implied]. Below, we summarize the approaches that are relevant to our methodology.
The simplest way to obtain the function $\hat{\sigma}(T,K)$ is to choose an interpolation and extrapolation scheme to interpolate the available volatility quotes on the continuum. It is well known [fengler2009arbitrage] that such an approach is not arbitrage-free since a direct interpolation on the volatility surface often leads to a non-convex option pricing function. The mispricing is further exaggerated for assets with few option quotes available in the market, where the interpolation determines a large portion of the surface, increasing the risk of generating implausible or arbitrageable prices. An arbitrage-free interpolation scheme is feasible on the option pricing function rather than the implied volatilities [le2021arbitrage, andreasen2011volatility]. The interpolation methods remain problematic for assets for which only a few high-quality option prices are available. Furthermore, to obtain the implied volatilities from the market prices, an inversion of the Black-Scholes formula is required, which can only be achieved by a root-finding algorithm.
@@
Volatility parametrizations are often directly derived from a parametric asset model, such as a stochastic volatility model. Although a volatility surface can be generally derived for any asset price model, not all models offer a convenient analytical form for the volatility surface. Defining an analytically tractable and realistic model (in terms of fitting the market) is thus a challenging task. The Stochastic Volatility Inspired (SVI) parametrization [gatheral2014arbitrage] was derived from a stochastic volatility model and enjoyed great success in the past. The parametrization was extended and improved in multiple works [mingone2022no, corbetta2019robust]. Another successful parametrization is the SABR parametrization [Hagan:2002], a direct result of the Hagan et al. formula provided by the SABR model (Stochastic-Alpha-Beta-Rho). Parametric volatility surfaces are convenient for numerous reasons. Firstly, since the parametrizations are derived from an asset model, the absence of arbitrage is guaranteed as long as the parameters remain in a well-defined range. Simple constraints on the parameter space are sufficient to ensure the absence of arbitrage. Secondly, the parameters can be attributed to the shape of the volatility surface. This allows for a simple way to compare surfaces and the expression of the dynamics of the surface in terms of the dynamics of the parameter. Thirdly, theoretically, no minimum number of market quotes is required to fit an arbitrage-free surface. Although a higher number of quotes is desired for stability in the calibration, a parametrization can be fit to any number of quotes and remain arbitrage-free. This is of particular importance for assets with only a few quotes or only in a particular strike region, where little to no information is provided for the tails of the volatility curves.
@@
The main drawback of parametrizations is that there is no guarantee that a surface that is appropriately close to the market data can be found. The flexibility of the surface is limited by the chosen parametrization, which does not cover all viable (i.e. arbitrage-free) volatility shapes. If the market conditions are not within the scope of the parametrization, the fitting process fails, and the parametrization misrepresents the actual market conditions. These out-of-scope market conditions can be systemic, meaning that parametric models are generally unfit to model a certain market behavior. For instance, it is well-known that short-maturity options in equity markets exhibit a steeper at-the-money implied volatility term structure than attainable under regular stochastic volatility models [gatheral2011volatility]. Alternatively, the market condition can be out of scope due to irregular behavior in the market, such as spikes in volatility or higher-than-usual uncertainty. It is not uncommon for very short-term options markets (near expiry) to exhibit a W-shaped volatility shape, or “mustache” shapes [glasserman2023w] before an earnings announcement of the underlying equity. These shapes of volatility surface arise from bimodal risk-neutral probability density functions for the stock price, which reflect the dichotomous nature of the earnings, which can either have a positive or negative effect on the asset price. Classical diffusion models [^1] It needs to be mentioned that SVI-type parametrizations do not rely on the an underlying model, but their development was merely inspired by volatility shapes of the Heston model. Nevertheless, one observes the same characteristics in terms of fit as classical stochastic volatility models. are unable to create such shapes due to the single continuous diffusive driver, and the corresponding volatility parametrization is limited in such cases. This limitation highlights the industry’s need for more flexible parameterized surfaces to achieve better fits in such market situations.
@@ 1.3 Contributions
In this paper, we study volatility surfaces in cases when the parametric surfaces reach the limit of their flexibility. We introduce a generic method to enhance the flexibility in such cases using a randomization scheme of the parameters of the surface. By replacing one or more parameters of the parametric surface with a random variable, we derive a new implied volatility surface with increased flexibility to fit the market. The additional flexibility stems from an induced mixing probability density, which is created by the stochastic nature of the randomized parameter. Mixing techniques have been studied in the past and have been shown to fit well to the market [rebonato2004unconstrained, bloch2011smiling, wilkens2005option]. The new randomized volatility surface is expressed in terms of the original parameters plus parameters to specify the chosen random variable. To show the effectiveness of the novel method, we apply the method to the existing parametrizations of the SABR parametrization and demonstrate that the additional flexibility enables them to better fit the market data for SPX options with expiry dates of less than half a year. Furthermore, we derive a second kind of randomization where we randomize the spot parameter of the asset. This arbitrage-free randomization is shown to be particularly effective for near-maturity options (i.e., options with an expiry date in the next few days) before an earning announcement, which induces a multi-modal-type volatility regime. [^2] We speak of a multi-modal-type volatility regime when the risk-neutral probability density functions implied by the volatility surface exhibit more than one mode and are therefore multi-modal. The paper is organized in the following way: Section 2 describes the process of randomizing a parametrization of volatility surfaces and proves that the randomization is indeed free of static arbitrage. We present a few examples of randomized parametrizations and their effectiveness in Section 3. In particular, we show an improved fit to options on the S&P 500 index. In Section 4, we then introduce the randomized spot parametrization and also present an example of its use on short-maturity AMZM (Amazon.com Inc) options on the day of an earnings announcement. Finally, we conclude the paper in Section 5.
@@ 2.1 General Volatility Parametrizations
We consider a type of implied volatility surfaces, which are given in parametric form, defined by a set of parameters. As a starting point, suppose that $\overline{p}:=(p_{1},\dots,p_{m})$ is an array of $m$ constant parameters on a parameter domain $\overline{p}\in\mathcal{D}\subset\mathbb{R}^{m}$. We define a parameterized implied volatility function as a positive $\mathcal{C}^{1,2}$ function
@@
for an asset with spot price $S_{0}$ at $t_{0}$. The function provides a continuum of European call and put option prices on the underlying asset on $(t_{0}\times\infty)\times[0,\infty)$ given the implied volatility $\hat{\sigma}(T,K;\overline{p})$. We can extend the functions to the limit points at $T=t_{0}$, which, under certain conditions exist and are given by $V_{c}(t_{0},K;\overline{p})=(S_{0}-K)^{+}$ and $V_{c}(t_{0},K;\overline{p})=(K-S_{0})^{+}$ (see 2.1). The pricing function defines two additional surfaces in $\mathbb{R}^{3}$ over the domain $\Pi$, which we refer to as the (put/call) pricing surface.
Since the parameterized implied volatility and pricing surface define a set of prices for tradable options, these prices must be free of arbitrage opportunities. There are various ways to define the absence of arbitrage of an implied volatility surface, most of which are equivalent. Here, we will utilize the model-free definition [roper2009implied], which does not rely on the introduction of stochastic asset models but can be expressed as a set of conditions on the pricing surfaces.
@@
Equivalent conditions in terms of the put-call parity can be derived for the put pricing function $V_{p}(T,K;\overline{p})$, or directly in terms of the volatility surface $\hat{\sigma}(T,K;\overline{p})$ [guo2016generalized]. Under these conditions, one can prove the existence of a non-negative local martingale process on a suitable probability space such that the call price function can be written as a risk-neutral expectation of the final payoff. In other words, an arbitrage-free model (price process) exists that yields the volatility surface. However, we will not utilize this fact to avoid introducing any model dynamics and remain “model-free”.
Due to the continuity of the Black-Scholes formula with respect to the volatility input, and since we chose the volatility surface to be $\mathcal{C}^{1,2}$, it is necessary for the parameterized price function $V_{c/p}(T,K;\overline{p})$ to be $\mathcal{C}^{1,2}$. Therefore, one can obtain the risk-neutral probability density functions $\{p_{S_{t}}\colon t\geq t_{0}\}$, defined as [^4] Note that one can also use the put prices $V_{p}(t,x;\overline{p})$ to obtain the same result.
@@
for any $t\geq t_{0}$ using the Breeden-Litzenberger formula [breeden1978prices] and risk-free rate $r$. In particular, one can show with a few manipulations that the expected value of the random variable $S_{t}$ whose PDF is given by is $p_{S_{t}}(\cdot)$, is
@@
Conversely, given a set of PDFs, such that (2.4) holds for all $t\geq t_{0}$, we can define an arbitrage-free pricing surface by integration of the probability densities [gatheral2014arbitrage].
@@
with respect to a desired norm $\lVert.\rVert$. The ability to fit the market, therefore, directly depends on the “size” of $\mathcal{S}$, whether $\mathcal{S}$ contains a function that can match the market.
@@ 2.2 Volatility Parametrizations with Random Parameters
In an effort to increase the size of $\mathcal{S}$ and obtain better fitting arbitrage-free volatility surfaces, we consider the possibility of adding stochasticity to the implied volatility surface by replacing one of the parameters $p_{i}$ for $1\leq i\leq m$ with a random variable. Let $\hat{\sigma}(T,K;\overline{p})$ be a parameterized implied volatility function with parameter domain $\mathcal{D}\subset\mathbb{R}^{m}$ and suppose that $\vartheta$ is a real-valued random variable on a probability space $(\Omega,\mathcal{F},\mathbb{P})$. We can replace the $i$-th entry of $\overline{p}$ with the random variable $\vartheta$ to obtain the random vector $\overline{p}(\vartheta)=(p_{1},p_{2},\dots,\vartheta,\dots,p_{m})$. If we assume that $\overline{p}(\vartheta)$ is almost surely in the domain $\mathcal{D}$, a realization $\omega\in\Omega$ determines a real-valued vector $\overline{p}(\theta):=(p_{1},p_{2},\dots,\theta,\dots,p_{m})\in\mathbb{R}^{m}$ with $\theta=\vartheta(\omega)$, which almost surely provides an arbitrage-free implied volatility surface $\hat{\sigma}(T,K;\overline{p}(\theta))$ and its associated pricing functions $V_{c/p}(T,K;\overline{p}(\theta))$.
Assuming measurability of the parametrization function in the parameters, the pricing function $V_{c/p}(T,K;\overline{p}(\vartheta))$ is thus a random variable, and we can compute the expectation as the Lebesgue-integral
$\mathbb{E}[V_{c/p}(T,K;\overline{p}(\vartheta))]=\int BS_{c/p}(t_{0},S_{0},T,K;\hat{\sigma}(T,K;\overline{p}(\vartheta))){\rm d}\mathbb{P}.$ (2.7)
If $\vartheta$ is absolutely continuous, the integral is a Riemann integral
@@
where $f_{\vartheta}$ is the associated probability density function of $\vartheta$. The equation shows that the expected option price under the randomized volatility surface is effectively an average of option prices on the sample space of the parameters of the volatility surface. Since the expectation of the randomized pricing function is a deterministic function in two variables, we can investigate its suitability as a pricing function. Due to the linear properties of the integral operator, the function is indeed an arbitrage-free pricing function.
@@
Let $\hat{\sigma}(T,K;\overline{p})$ be an arbitrage-free parameterized implied volatility function on the parameter domain $\overline{p}\in\mathcal{D}$ and let $\vartheta$ be a real-valued random variable on a probability space $(\Omega,\mathcal{F},\mathbb{P})$ and denote $\mu(\cdot)$ as its law.
Suppose that $\overline{p}(\vartheta)=(p_{1},p_{2},\dots,\vartheta,\dots,p_{m})$ is the random vector where we replaced the $i$-th parameter of $\overline{p}$ with $\vartheta$. Then, the map $\hat{\sigma}(T,K)$ such that
@@
$V_{c}(T,K;\overline{p}(\theta))={\rm e}^{-r(T-t_{0})}\int_{K}^{\infty}(x-K)p_{S_{T};\theta}(x){\rm d}x.$
@@
$\mathbb{E}[V_{c}(T,K;\overline{p}(\vartheta))]$ $=\int_{\mathbb{R}}V_{c}(T,K;\overline{p}(\theta)){\rm d}\mu(\theta)$
$=\int_{\mathbb{R}}{\rm e}^{-r(T-t_{0})}\int_{K}^{\infty}(x-K)p_{S_{T};\theta}(x){\rm d}x{\rm d}\mu(\theta)$
$={\rm e}^{-r(T-t_{0})}\int_{K}^{\infty}(x-K)\int_{\mathbb{R}}p_{S_{T};\theta}(x){\rm d}\mu(\theta){\rm d}x.$
Since the random vector $\overline{p}(\vartheta)$ takes values almost surely in $\mathcal{D}$, the function $p_{S_{T};\vartheta}(\cdot)$ is almost surely a probability density function. The expression $\int_{\mathbb{R}}p_{S_{T};\theta}(x){\rm d}\mu(\theta)$ is a convex combination of probability densities (mixture density), which we define as the function
$\overline{f}(x):=\int_{\mathbb{R}}p_{S_{T};\theta}(x){\rm d}\mu(\theta).$
We claim $\overline{f}(x)$ to be a proper probability density function on $0,\infty)$. The function $\overline{f}(x)$ is certainly positive since $p_{S_{T};\theta}(\cdot)$ is a positive function, and integrates to 1 since
$\int_{0}^{\infty}\overline{f}(x){\rm d}x$ $=\int_{0}^{\infty}\int_{\mathbb{R}}p_{S_{T};\theta}(x){\rm d}\mu(\theta){\rm d}x$
$=\int_{\mathbb{R}}\int_{0}^{\infty}\ p_{S_{T};\theta}(x){\rm d}x{\rm d}\mu(\theta)$
$=\int_{\mathbb{R}}{\rm d}\mu(\theta)$
@@
This show the sufficient condition for $\hat{\sigma}_{T}(K)$ to be butterfly-arbitrage-free [gatheral2014arbitrage]. It remains to show that condition v) is true as well. Let $K$ be fixed and suppose that $t_{1}<t_{2}\in\Pi_{T}$. Since $\hat{\sigma}(T,K;\overline{p}(\theta))$ is arbitrage-free, we have that $V_{c}(t_{1},K;\overline{p}(\theta))\leq V_{c}(t_{2},K;\overline{p}(\theta))$ for $\overline{p}(\theta)\in\mathcal{D}$. Since this is almost surely the case, we conclude that